# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

### v1

@@ 3 Curvature assumptions and operator analysis
We start by adding assumptions on (1) for use in later subsections. Given $h:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ and $f_{i}:\operatorname*{dom}h\mapsto\mathbb{R}$ for $i\in[k]$, define the assumptions:
@@
We now give a few remarks about (A1)–(A2). First, (A1) is necessary to ensure that $\operatorname{prox}_{h}(\cdot)$ is well-defined. Second, it can be shown [^4] See, for example, [5, 11] and [10, Proposition 2.1.55]. that (A2) is equivalent to the assumption that $\nabla f$ is $\max\{m,M\}$-Lipschitz continuous when $m=M$. Third, the lower bound in (9) is equivalent to assuming that $f_{i}(\cdot)+m\|\cdot\|^{2}/2$ is convex and, hence, if $m=0$ then $f_{i}$ is convex. The parameter $m$ is often called a weak-convexity parameter of $f_{i}$. Fourth, using symmetry arguments and the third remark, if $M=0$ then $f_{i}$ is concave. Finally, the third remark motivates why we choose two parameters, $m$ and $M$, in (9). Specifically, we use $m$ (resp. $M$) to develop results that can be described in terms of the level of convexity (resp. concavity) of the problem.
@@
Clearly, when $\operatorname{prox}_{q_{i}}(y)=y$ for every $y\in\mathbb{R}^{n}$, the above update corresponds to a SGD step applied to $\min_{x}\sum_{i=1}^{k}f_{i}(x)$ with stepsize $\lambda$ and starting point $x$. Moreover, while it is straightforward to show that ${\cal A}_{\lambda}(\cdot)$ is $(1+\lambda\max\{m,M\})$-Lipschitz continuous when $\{f_{i}\}$ satisfies (A2) [^5] This follows by using the triangle inequality and the non-expansiveness of $\operatorname{prox}_{q_{i}}(\cdot)$., we prove that the Lipschitz constant can be improved to $\Theta(\sqrt{1+\lambda m})$ when $\lambda$ is small. Notice, in particular, that the former bound does not converge to one when $m\to 0$, e.g., when $f_{i}$ becomes more convex, while the latter does.
To start, we recall the following well-known bound from convex analysis. Its proof can be found, for example, in [5, Theorem 5.8(iv)].
Lemma 3.1.
if and only if
Lemma 3.2.
$\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle\geq-\left[m+\frac{1}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.$
Proof.
Define $F=f_{i}+m\|\cdot\|^{2}/2$ and let $x,y\in\operatorname*{dom}h$ be fixed. Moreover, note that $F$ is convex and satisfies (11) with $L=M+m$. It then follows from Lemma 3.1 with $L=M+m$ that
$\leq\left\langle\nabla F(x)-\nabla F(y),x-y\right\rangle$
$\geq-\left[m+\frac{m^{2}}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.\qed$
Lemma 3.3.
$\Delta:=u-v,\quad\Delta^{p}:=\operatorname{prox}_{\psi}(u)-\operatorname{prox}_{\psi}(v).$
Then, the following statements hold:
  • (a)
    $\|\Delta^{p}\|\leq\left\langle\Delta,\Delta^{p}\right\rangle$;
  • (b)
    $\|\Delta^{p}-\Delta\|^{2}\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}$.
Proof.
(a) See [5, Theorem 6.42(a)].
(b) Using part (a), we have that
$\|\Delta^{p}-\Delta\|^{2}$ $=\|\Delta^{p}\|+\|\Delta\|^{2}-2\left\langle\Delta,\Delta^{p}\right\rangle\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}.\qed$
We are now ready to present the crucial properties of ${\cal A}_{\lambda}(\cdot)$.
@@
Define the scalar
$L_{\lambda}=L_{\lambda}(m,M):=\sqrt{1+2\lambda m\left[1+\frac{m}{M+m}\right]}\quad\forall\lambda>0.$ (12)
@@
    if $f_{i}$ satisfies (A2) and $\lambda\leq 1/[2(M+m)]$, then ${\cal A}_{\lambda}(\cdot)$ is $L_{\lambda}$-Lipschitz continuous.
Proof.
$\leq\|x-y\|+2\lambda C.$
(b) Let $x,y,z\in\operatorname*{dom}h$ be arbitrary, and suppose $\lambda\leq 1/(M+m)$. Moreover, denote
$\begin{gathered}\xi(\cdot):={\cal A}_{\lambda}(\cdot,\{f_{i}\},\{q_{i}\}),\quad d:=x-y,\\ \Delta_{i}:=\nabla f_{i}(x)-\nabla f_{i}(y),\quad\Delta_{i}^{p}:=\operatorname{prox}_{q_{i}}(\nabla f_{i}(x))-\operatorname{prox}_{q_{i}}(\nabla f_{i}(y)).\end{gathered}$
Using the previous bound, the fact that $\|\sum_{i=1}^{d}v_{i}\|^{2}\leq d\sum_{i=1}^{d}\|v_{i}\|^{2}$ for any $\{v_{i}\}\subseteq\mathbb{R}^{n}$, and Lemma 3.3(b) with $(\Delta,\Delta^{p})=(\Delta_{i},\Delta_{i}^{p})$, we have that
$=\frac{1}{|B|^{2}}\left\|\sum_{i\in B}(d-\lambda\Delta_{i}^{p})\right\|^{2}\leq\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}$
$=\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda(\Delta_{i}+\Delta_{i}-\Delta_{i}^{p})\right\|^{2}=\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda\|\Delta_{i}-\Delta_{i}^{C}\|^{2}\right)$
$=\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda^{2}\|\Delta_{i}\|^{2}-\lambda^{2}\|\Delta_{i}\|^{2}\right)\leq\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda^{2}\|\Delta_{i}\|^{2}\right)$
$=2\left[\|d\|^{2}+\lambda\left(2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\right)\right].$ (14)
Let us now bound the last two terms of the above expression. Using Lemma 3.1 for the second term, and the fact that $\lambda\leq 1/[2(M+m)]$, we have
$2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\leq 2\lambda\sum_{i\in B}\|\Delta_{i}\|^{2}+2\left[m+\frac{m^{2}}{2(M+m)}\right]\|d\|^{2}-\frac{\sum_{i\in B}\|\Delta_{i}\|^{2}}{(M+m)|B|}$
$=2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}+\left(\underbrace{2\lambda-\frac{1}{M+m}}_{\leq 0}\right)\sum_{i\in B}\|\Delta_{i}\|^{2}\leq 2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}.$ (15)
Using (14), (15), and the definition of $d$, we conclude that
$\|\xi(x)-\xi(y)\|^{2}$ $\leq\|d\|^{2}+\lambda\left(2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\right)$
$\leq\|d\|^{2}+2\lambda m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}=L_{\lambda}^{2}\|x-y\|^{2}.\qed$
@@
Before ending this subsection, we give a few remarks about part (b) of the above result. First, $L_{\lambda}(0,M)=1$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive when $f_{i}$ is convex for every $i\in[k]$ and $\lambda\leq 1/(2M)$. Second, if $\lambda=1/(2m)$ then $L_{\lambda}(m,0)=\sqrt{3}$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ $\sqrt{3}$-Lispchitz continuous when $f_{i}$ is concave. Third, like the first remark, $L_{0}(m,M)=1$ implies that ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive. Finally, when $m=M$ and $\lambda=1/(2m)$, we have $L_{\lambda}(m,M)=\sqrt{5/2}$ and ${\cal{A}}_{\lambda}(\cdot)$ is $\sqrt{5/2}$-Lispchitz continuous.
4 Privacy bounds for DP-SGD
This section develops RDP bounds for the last iterates of DP-SGD under cyclic batch sampling.
For ease of reference, we define
$q_{C}(x)=\begin{cases}1,&\text{if }\|x\|\leq C,\\ \infty,&\text{otherwise},\end{cases}\quad{\rm Clip}_{C}(y)=\operatorname{prox}_{q_{C}}(y).$ (16)
and formally present DP-SGD in Algorithm 1.
Algorithm 1 Cyclically-sampled last-iterate DP-SGD
1:Inputs: $\{f_{i}\}_{i=1}^{k}$, $h$, samples $\{N_{t}\}\subseteq\mathbb{R}^{n}$ from ${\cal N}(0,\sigma^{2}I)$, and $X_{0}\in\operatorname*{dom}h$;
2:Parameters: batch size $b$, stepsize $\lambda$, clipping norm $C$, and iteration limit $T$;
3:Outputs: $X_{T}\in\operatorname*{dom}h$;
4:for $t\leftarrow 1,\ldots,T$ do
5:  $j_{t}\leftarrow b(t-1)\bmod k$
6:  $B_{t}\leftarrow\{j_{t}+1,\ldots,j_{t}+b\}$
7:  $g_{t}\leftarrow(1/b)\sum_{i\in B_{t}}{\rm Clip}_{C}(\nabla f_{i}(X_{t-1}))$
8:  $X_{t}\leftarrow\operatorname{prox}_{\lambda h}\left(X_{t-1}-\lambda g_{t}+N_{t}\right)$
9:return $X_{T}$
Before proceeding, let us give a few remarks about Algorithm 1. First, $j_{t}$ corresponds to the first index of the batch indices $B_{t}$, and if $j_{t}>j_{t-1}$ then $j_{t}-j_{t-1}=b$. Second, $g_{t}$ corresponds to the average of the clipped gradients of $f_{i}$ at $X_{t-1}$. Third, since each $X_{t}$ is generated by a call to $\operatorname{prox}_{\lambda h}(\cdot)$, it holds that $X_{t}\in\operatorname*{dom}h$ for every $t\in[T]$. Finally, notice that the noise $N_{t}$ is added to the displacement $X_{t-1}-\lambda g_{t}$ rather than to $g_{t}$. This setup allows us to develop RDP bounds that scale with both $\lambda$ and $\sigma$.
@@ 4 Privacy bounds for DP-SGD
Now, for the remainder of this section, suppose $h$ satisfies (A1) and let $f_{i}^{\prime}:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ for $i\in[k]$ be such that there exists $i^{*}\in[k]$ where $f_{i}^{\prime}=f_{i}$ for every $i\neq i^{*}$ and $f_{i^{*}}^{\prime}\neq f_{i^{*}}$, i.e., $\{f_{i}\}\sim\{f_{i}^{\prime}\}$. In other words, $i^{*}$ is the index where the neighboring datasets $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ differ.
For later reference, we also define the assumption:
@@
The next result shows how the updates in Algorithm 1 are randomized proximal updates applied to the operator $A_{\lambda}(\cdot)$ in (10) with $q_{i}(\cdot)=q_{C}(\cdot)$ from (16).
Lemma 4.1.
Let $\{X_{t}\}$ and $\{X_{t}^{\prime}\}$ denote the iterates generated by Algorithm 1 with functions $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$, respectively, and fixed $\lambda$, $C$, $b$, $\sigma$, $\{N_{t}\}$, $T$, and $X_{0}$ for both sequences of iterates. Moreover, denote
$\phi_{t}(x):={\cal A}_{\lambda}(x,\{f_{i}\},\{q_{C}\}),\quad\phi_{t}^{\prime}(x):={\cal A}_{\lambda}(x,\{f_{i}^{\prime}\},\{q_{C}\}),\quad\forall x\in\operatorname*{dom}h,$
where $q_{C}$ and ${\cal A}_{\lambda}$ are as in (16) and (10), respectively. Then, it holds that
Proof.
This follows immediately from the definition of $\phi_{t}$ and the update rules in Algorithm 1. ∎
We now present some important norm bounds.
Lemma 4.2.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\{\phi_{t}\}$, and $\{\phi_{t}^{\prime}\}$ be as in Lemma 4.1 and denote $\ell=k/b$ and $t^{*}:=\inf_{t\geq 1}\left\{t:i^{*}\in B_{t}\right\}$. Then, it holds that
Proof.
$\leq\frac{\lambda}{b}\left[\|{\rm Clip}_{C}(\nabla f_{i^{*}}(x))\|+\|{\rm Clip}_{C}(\nabla f_{i^{*}}^{\prime}(x))\|\right]=\frac{2\lambda C}{b}.\qed$
In the remaining subsections, we show how to use the accumulated results (primarily, Proposition 2.2) to obtain the desired RDP bounds.
4.1 Curvature-independent bounds
This section establishes an RDP bound without assuming any additional conditions on $\{f_{i}\}$. Specifically, using Proposition 2.2(b), we obtain the following bound.
@@ 4.1 Curvature-independent bounds
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $\sigma$, $C$, $\ell$ and $T$ be as in Lemmas 4.1 and 4.2. If $T\geq\ell$, then for any $\lambda>0$ it holds that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 8\alpha T\left(\frac{\lambda C}{\sigma}\right)^{2}.$ (18)
Proof.
Remark that if $Y_{0}=X_{t^{*}}$, $Y_{0}^{\prime}=X_{t^{*}}^{\prime}$, and $\ell=k/b$, then Proposition 3.4(a), Lemma 4.2, and the fact that $b\geq 1$ imply that the assumptions of Proposition 2.2(b) hold with $\zeta=\bar{s}=2\lambda C$. Consequently, using Proposition 2.2(b) with the previous substitutions, $X_{T}=Y_{T-t^{*}-1}$, $X_{T}^{\prime}=Y_{T-t^{*}-1}^{\prime}$, and $T=T-t^{*}-1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})=D_{\alpha}(Y_{T-t^{*}-1}\|Y_{T-t^{*}-1}^{\prime})\leq 8\alpha(T-t^{*}-1)\left(\frac{\lambda C}{\sigma}\right)^{2}\leq 8\alpha T\left(\frac{\lambda C}{\sigma}\right)^{2}.\qed$
We now give two remarks. First, the right-hand-side of the above bound can be made small by decreasing $\lambda$ or increasing $\sigma$. [^6] While decreasing $\lambda$ generally reduces the amount of Gaussian noise added, it can be detrimental to the convergence of the model. This highlights the inherent privacy/accuracy trade-off of DP-SGD [4]. However, the upper bound cannot be made small by increasing the batch size $b$. Second, it improves on the naive bound that is obtained from bounding $X_{T}$ using the triangle inequality and the clipping step. More specifically, using the update rule for $X_{t}$ and the triangle inequality, it is straightforward to show that $\|X_{T}\|\leq\|X_{0}\|+\lambda TC$. Hence, using the Rényi divergence bound between two Gaussian random variables that differ in their mean parameter, one can replace the right-hand-side of (18) with a $\Theta(\alpha[\lambda CT/\sigma]^{2})$ term, which is a factor $T$ worse than the bound in (18).
In the next subsection, we develop improved RDP upper bounds that decrease with increasing $b$, under more restrictive assumptions on $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$.
4.2 Curvature-dependent bounds
This section establishes two RDP bounds under the assumption that $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3) and $\lambda$ is small relative to $1/(m+M)$.
For the first bound, suppose $\operatorname*{dom}h$ is bounded with diameter
$d_{h}:=\sup\{\|x-y\|:x,y\in\operatorname*{dom}h\}<\infty.$ (19)
Using a special application of Proposition 2.2(a), we obtain the following $d_{h}$-dependent bound.
Theorem 4.4.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $\sigma$, $C$, and $T$ be as in Lemma 4.1. If $\lambda\leq 1/[2(m+M)]$ and $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3), then
@@ 4.2 Curvature-dependent bounds
Proof.
Using Proposition 2.2(a) with
and $E=T=1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq D_{\alpha}^{(d_{h})}(X_{T-1}\|X_{T-1}^{\prime})+\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}\theta_{L}(1)=\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}.\qed$
Using a special application of Proposition 2.2(b), we present a bound that is linear in terms of the number passes $E$ through the dataset. For conciseness, we leave its proof in Appendix A.3.
@@
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $C$, and $T$ be as in Lemma 4.1 and denote $E:=\lfloor T/\ell\rfloor$ and $\ell:=k/b$. If $\lambda\leq 1/[2(m+M)]$, $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3), and $T\geq\ell$, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[\theta_{L_{\lambda}}(T-E\ell)+E\cdot\theta_{L_{\lambda}}(\ell)\right],$ (21)
where $L_{\lambda}$ is as in (12). Moreover, if $f_{i}$ is convex for every $i\in[k]$, then $m=0$ and
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left(\max\left\{0,\frac{1}{T-E\ell}\right\}+\frac{E}{\ell}\right).$ (22)
We conclude with a few remarks about the above bounds. First, the bound in (22) is on the same order of magnitude as the bound in [8] in terms of $T$ and $\ell$. However, the right-hand-side of (22) scales linearly with a $\lambda^{2}$ term, which does not appear in [8]. Second, as $\theta_{L_{\lambda}}(\cdot)\leq 1$, the right-hand-side of (21) increases (at most) linearly with respect to the number of dataset passes $E$. Third, substituting $\sigma=\Theta({C/[b\sqrt{\epsilon}]})$ in (21) yields a bound that depends linearly on $\varepsilon$ and i invariant to changes in $C$ and $b$.
In Appendix C, we discuss further choices of the parameters in (21) and their implications. Some highlights are as follows. First, when $\ell$ is sufficiently large, the terms in (21) that involve $\theta_{L_{\lambda}}(\ell)$ become $\Theta(1+Em/[M+m])$, which shows that the effect of $E$ diminishes as $m\to 0$ (with the caveat that $\ell$ grows as well). Second, when $\lambda\leq\min\{1/\sqrt{E},1/2[m+M]\}$, the resulting bound is independent of $E$.
Appendix A Technical proofs
A.1 Proof of Lemma 2.1
The below result gives two calculus rules for the shifted Rényi divergence given in (3). In particular, the proof of the second rule is a minor modification of the proof given for [8, Lemma 21].

### v3

@@ 2 Privacy bounds for DP-SGD
We start by precisely giving the assumptions on (1). Given $h:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ and $f_{i}:\operatorname*{dom}h\mapsto\mathbb{R}$ for $i\in[k]$, consider the following assumptions:
@@
We now give a few remarks about (A1)–(A2). First, (A1) is necessary to ensure that $\operatorname{prox}_{h}(\cdot)$ is well-defined. Second, it can be shown [^14] See, for example, [7, 26] and [21, Proposition 2.1.55]. that (A2) is equivalent to the assumption that $\nabla f$ is $M$-Lipschitz continuous when $m=M$. Third, the lower bound in (11) is equivalent to assuming that $f_{i}(\cdot)+m\|\cdot\|^{2}/2$ is convex and, hence, if $m=0$ then $f_{i}$ is convex. The parameter $m$ is often called a weak-convexity parameter of $f_{i}$. Fourth, using symmetry arguments and the third remark, if $M=0$ then $f_{i}$ is concave. Finally, the third remark motivates why we choose two parameters, $m$ and $M$, in (11). Specifically, we use $m$ (resp. $M$) to develop results that can be described in terms of the level of convexity (resp. concavity) of the problem.
@@
Clearly, when $\operatorname{prox}_{q_{i}}(y)=y$ for every $y\in\mathbb{R}^{n}$, the above update corresponds to a SGD step applied to the problem of minimizing $\sum_{i=1}^{k}f_{i}(z)$ (with respect to $z$) under the stepsize $\lambda$ and starting point $x$. Moreover, while it is straightforward to show that ${\cal A}_{\lambda}(\cdot)$ is $(1+\lambda\max\{m,M\})$-Lipschitz continuous when $\{f_{i}\}$ satisfies (A2) [^15] See, for example, [11, Appendix A.6]., we prove in Proposition 2.1(b) that the Lipschitz constant can be improved to $\sqrt{1+\kappa\lambda m}$ for some $\kappa\leq 4$ when $\lambda$ is small. Notice that the former constant does not converge to one when $m\to 0$, e.g., when $f_{i}$ becomes more convex, while the latter one does.
We are now present some important properties of ${\cal A}_{\lambda}(\cdot)$.
@@
Let $(m,M)$ be as in assumption (A2), and define
$L_{\lambda}=L_{\lambda}(m,M):=\sqrt{1+2\lambda m\left[1+\frac{m}{2(M+m)}\right]}\quad\forall\lambda>0.$ (13)
@@
    if $\{f_{i}\}$ satisfies (A2) and $\lambda\leq 1/[2(M+m)]$ then ${\cal A}_{\lambda}(\cdot)$ is $\sqrt{2}L_{\lambda}$-Lipschitz continuous;
  • (c)
@@
    if $\{f_{i}\}$ satisfies (A2), $\lambda\leq 1/(M+m)$, and $\nabla f_{i}(x)=\operatorname{prox}_{q_{i}}(\nabla f_{i}(x))$ for every $i\in[k]$ and $x\in\operatorname*{dom}h$, then ${\cal A}_{\lambda}(\cdot)$ is $L_{\lambda}$-Lipschitz continuous on $\operatorname*{dom}h$.
@@
Some remarks are in order. First, $L_{\lambda}(0,M)=1$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive when $f_{i}$ is convex for every $i\in[k]$, $\lambda\leq 1/M$, and $\nabla f(x_{i})=\operatorname{prox}_{q_{i}}(\nabla f(x_{i}))$. Second, if $\lambda=1/(2m)$ then $L_{\lambda}(m,0)=\sqrt{3}$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ $\sqrt{6}$-Lispchitz continuous when $f_{i}$ is concave. Third, like the first remark, $L_{0}(m,M)=1$ implies that ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive. Finally, when $m=M$ and $\lambda=1/(2m)$, we have $L_{\lambda}(m,M)=\sqrt{5/2}$ and ${\cal{A}}_{\lambda}(\cdot)$ is $\sqrt{5}$-Lispchitz continuous.
@@
For the remainder of this section, suppose $h$ satisfies (A1) and let $f_{i}^{\prime}:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ for $i\in[k]$ be such that there exists $i^{*}\in[k]$ where $f_{i}^{\prime}=f_{i}$ for every $i\neq i^{*}$ and $f_{i^{*}}^{\prime}\neq f_{i^{*}}$, i.e., $\{f_{i}\}\sim\{f_{i}^{\prime}\}$. That is, $i^{*}$ is the index where the neighboring datasets $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ differ.
We also make use of the follow assumption.
@@
We now present the main RDP bounds in terms of the following constants:
$d_{h}:=\sup\{\|x-y\|:x,y\in\operatorname*{dom}h\},\quad\theta_{L}(0):=0,\quad\theta_{L}(s):=\frac{L^{2(s-1)}}{\sum_{j=0}^{s-1}L^{2j}}\quad\forall s\geq 1.$ (15)
We first present a bound where $\operatorname*{dom}h$ is bounded with diameter $d_{h}<\infty$.
@@
Let $\{X_{t}\}$ and $\{X_{t}^{\prime}\}$ denote the iterates generated by Algorithm 1 with per-example loss functions $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$, respectively, and fixed $\lambda$, $C$, $b$, $\sigma$, $\{N_{t}\}$, $T$, and $X_{0}$ for both sequences of iterates. If $\lambda\leq 1/[2(m+M)]$ and (A1)–(A3) hold, then
@@
We now present the RDP bounds for when $\operatorname*{dom}h$ is (possibly) unbounded.
@@
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $\sigma$, $b$, $C$, and $T$ be as in Theorem 2.2, and denote $\ell:=k/b$ and $E:=\lfloor T/\ell\rfloor$. If $\lambda\leq 1/[2(m+M)]$ and (A1)–(A3) hold, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[1+E\cdot\theta_{\sqrt{2}L_{\lambda}}(\ell)\right],$ (17)
where $L_{\lambda}$ and $\theta_{L}(\cdot)$ are as in (13) and (15), respectively. On the other hand, if
$\max_{i\in[k],t\in[T]}\{\|\nabla f_{i}(X_{t})\|,\|\nabla f_{i}(X_{t}^{\prime})\|\}\leq C,$ (18)
i.e., no gradient clipping is performed, and $\lambda\leq{1}/{(M+m)}$ then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[1+E\cdot\theta_{L_{\lambda}}(\ell)\right].$ (19)
We conclude with a few remarks about the above bounds. First, the bound in (19) is on the same order of magnitude as the bound in [17] in terms of $T$ and $\ell$ when $L_{\lambda}=1$. However, the right-hand-side of (19) scales linearly with a $\lambda^{2}$ term, which does not appear in [17]. Second, as $\theta_{L_{\lambda}}(\cdot)\leq 1$, the right-hand-sides of (17) and (19) increases (at most) linearly with respect to the number of dataset passes $E$. Third, substituting $\sigma=\Theta({C/[b\sqrt{\epsilon}]})$ in (17) yields a bound that depends linearly on $\varepsilon$ and is invariant to changes in $C$ and $b$. In Appendix C, we discuss further choices of the parameters in (19) and their properties.
Appendix A Derivation of main results
This appendix derives the main results, namely, Theorems 2.2 and 2.3. It contains three subappendices. The first one derives important properties of a family of randomized operators, the second specializes these results to the DP-SGD update operator in (12), and the last one gives the proofs of Theorems 2.2 and 2.3 using the previous two subappendices.
A.1 General operator analysis
This subappendix gives some crucial properties about randomized proximal Lipschitz operators, which consist of evaluating a Lipschitz proximal operator followed by adding Gaussian noise. More specifically, it establishes several RDP bounds based on the closeness of neighboring operators.
We first bound the shifted Rényi divergence of a randomized proximal Lipschitz operator. The proof of this result is a straightforward extension of the argument in [17, Theorem 22] from $1$-Lipschitz operators to $L$-Lipschitz operators with additive residuals.
To begin, we present two calculus rules for the shifted Rényi divergence given in (5). In particular, the proof of the second rule is a minor modification of the proof given for [17, Lemma 21].